# Reweight-annealing method for calculating the value of partition function via quantum Monte Carlo

### v3

@@ Abstract
1D spin-1/2 Heisenberg chain.- As the first example, we consider an 1D spin-1/2 antiferromagnetic (AFM) Heisenberg chain with periodic boundary condition (PBC) with $\beta=30$ and length $L$. We use both the T-Re-An method and the Q-Re-An method to calculate the PF. For the quantum version, we write
where $H_{0}=\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$ and $H_{1}=\sum_{i\in\text{even}}{\mathbf{S}}_{i}\cdot{\mathbf{S}}_{i+1}$. Apparently, the unique ground state of $H_{1}$ is a tensor product of $L/2$ singlets, each carrying energy $-3/4$ ($\hbar=1$). The estimator for $Z(s^{\prime})/Z(s^{\prime\prime})$ is $\langle(s^{\prime}/s^{\prime\prime})^{n_{\text{odd}}}\rangle_{s^{\prime\prime}}$ in this model, with $n_{\text{odd}}$ denoting the amount of non-identity operators acting on all the bonds with odd indices within the framework of SSE. As it is shown in Fig. 2, both the results of T-Re-An and Q-Re-An method can match that of the exact diagonalization (ED) method exceptionally well for small lengths, which verify the validity of our methods. The results are also consistent with the expectation that $\ln Z\propto L+o(L)$, where $o(L)$ means higher order terms which is negligible compared with $L$. One may feel this example too trivial and easy to be calculated. Therefore we introduce the second example which needs a high-precision to extract the correction of the sub-leading term in entropy.
@@
2D LBW-HM.- To further exploit the advantage of our method, we calculate the thermodynamic entropy of a 2D lattice Bisognano-Wichmann Heisenberg model (LBW-HM) Giudici et al. (2018); Dalmonte et al. (2018), the thermodynamic entropy $S$ of the which at $\beta=1$ provides a good approximation for the Von Neumann entanglement entropy of a 2D PBC square lattice spin-1/2 AFM Heisenberg model under smooth bipartition Mendes-Santos et al. (2019, 2020). According to the CFT prediction and previous numerical results Metlitski and Grover (2015); Laflorencie (2016); Zhao et al. (2022a); Deng et al. (2023), the coefficient of sub-leading term of the entropy reflects the number of Goldstone modes. Here, the thermal entropy of the LBW-HM at $\beta=1$ should obey $S(L)=aL+b\ln L+c$, and $b=n_{G}/2$, where $n_{G}=2$ is the number of Goldstone modes of the 2D spin-1/2 AFM Heisenberg model Metlitski and Grover (2011); Laflorencie (2016). In this context, coarse result of entropy is disfavored because the the coefficient $b$ in front of the sub-leading term is sensitive to errors. We have compared our method with the specific heat integral ($C$-int) and Wang-Landau (WL) algorithm in this example.
@@
For the WL algorithm, on another hand, the allocation of total computational resources should not be predetermined as it hinges on when we can achieve a converged and flat histogram. Fig. 3 shows great error of the results obtained from the vanilla WL samplings Troyer et al. (2003). This reflects the problem of error saturation, which is intrinsic in this algorithm due to an approximately flat histogram and the empirical refinement parameter Belardinelli et al. (2008); Belardinelli and Pereyra (2007); Zhou and Bhatt (2005). Another disadvantage of the WL sampling compared with our method is that it cannot be parallelized: the two subroutines for flattening the corresponding histograms before and after an adjustment of the refinement parameter, must be executed sequentially. To achieve an arbitrary precision within the Wang-Landau framework is not easy. In contrast, our method is much more accessible as we only need to measure the observable of $Z({\mathbf{p}}_{k-1})/Z({\mathbf{p}}_{k})$ in many conventional and mature MC algorithms.
2D XXZ model.- As the last example, we consider a much more complex model, a 2D spin-1/2 XXZ model with next nearest neighbor interaction,
@@
For convenience, here we use equal distance of $T=1/\beta=0.01$ to implement our Re-An method, and it is also easy for numerical differentiation. Because a constant factor of the PF does not affect the (higher order) derivative of $F$, thus we choose $Z(T^{*}=1.95)$ as the reference to replace annealing to a known point. The PF we gained is actually $Z(T)/Z(T^{*})$ where $Z(T^{*})$ is an unknown constant. Then we indeed observe a divergent peak of the second order derivative of the free energy as shown in Fig. 4(a), which accurately probes the phase transition point, matching with the intersection point of Binder ratios $R_{2}=\langle M^{4}\rangle/\langle M^{2}\rangle^{2}$ of different sizes shown in Fig. 4(b), where $M$ denotes the magnetization.
@@
Summary.- We propose an unbiased algorithm within Monte Carlo framework, which has low-technical barrier and no systemic error. The method can be easily implemented for parallel computation and can obtain the value of partition function, free energy and thermal entropy with high precision. The method has two keynotes: (1) Find a parameter path in whole space (either temperature or external parameter) connecting a known point and the to-be-solved one. If we only want to calculate the (high order) derivative of the PF, even fixing an unknown point as the reference is enough; (2) Annealing along the path to calculate the final result through a propagation formula.
@@
We show the advantage of our method via calculating the thermal entropy of 2D LBW-HM and successfully extract precise coefficient of the sub-leading term ($\ln L$), which is consistent with the CFT prediction and catch the correct number of Goldstone modes. In comparison, the $C$-int and the WL sampling both have worse performance. In a 2D XXZ model, the numerical second derivative of the Re-An data is still smooth and precise, which accurately probes the phase transition point. Our method make the efficient and high-precision calculation of entropy and other related physical quantities accessible, which can reveal some universal information, such as CFT, criticality and symmetry breaking.
@@
Acknowledgement.- We acknowledge the start-up funding of Westlake University. The authors also acknowledge Beijing PARATERA Tech Co.,Ltd.(https://www.paratera.com/) for providing HPC resources that have contributed to the research results reported within this paper. CC and ZY thank the support of the National Natural Science Foundation of China (grant nos. 12174167, 12247101).
Appendix A Supplemental material
@@
A.1 (I) Estimators in thermal and quantum reweight-annealing under the SSE framework
For the thermal case, the estimator of $Z(\beta^{\prime})/Z(\beta^{\prime\prime})$ is general, which is
@@
$\begin{split}\frac{Z(\beta^{\prime})}{Z(\beta^{\prime\prime})}=&\frac{1}{Z(\beta^{\prime\prime})}\text{tr}\{e^{-\beta^{\prime}H}\}\\ =&\frac{1}{Z(\beta^{\prime\prime})}\text{tr}\bigg{\{}\sum_{n}\bigg{(}\frac{\beta^{\prime}}{\beta^{\prime\prime}}\bigg{)}^{n}\frac{(\beta^{\prime\prime})^{n}}{n!}(-H)^{n}\bigg{\}}\\ =&\bigg{\langle}\bigg{(}\frac{\beta^{\prime}}{\beta^{\prime\prime}}\bigg{)}^{n}\bigg{\rangle}_{\beta^{\prime\prime}}\\ \end{split}$ (1)
and the subscript denotes that this simulation should be performed at the inverse temperature $\beta^{\prime\prime}$.
@@
For the quantum case with $H(s)=sH_{0}+(1-s)H_{1}=H_{1}+s(H_{0}-H_{1})$, $Z(s^{\prime})/Z(s^{\prime\prime})$ usually takes the form $\langle(s^{\prime}/s^{\prime\prime})^{n_{s}}\rangle_{s^{\prime\prime}}$ within the framework of SSE, where $n_{s}$ is the number of non-identity operators related to $(H_{0}-H_{1})$. For instance, if
@@
where $H_{0}=\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$ and $H_{1}=\sum_{i\in\text{even}}{\mathbf{S}}_{i}\cdot{\mathbf{S}}_{i+1}$, then
@@
where $n_{\text{odd}}$ ($n_{\text{even}}$) denotes the amounts of non-identity operators acting on the bonds with odd (even) indices in SSE simulation, and
$\tilde{S}_{\alpha,M}=\frac{\beta^{n}(M-n)!}{M!}\prod_{i=1}^{M}\langle\alpha_{i-1}|H_{a_{i},b_{i}}|\alpha_{i}\rangle$ (4)
corresponds to operator string $S_{M}$ and the propagated states $\{\alpha_{0},\alpha_{1},\dots\alpha_{M}\}$ along the axis of imaginary time.
A.2 (II) Technical details of the annealing process
For the thermal and quantum reweight-annealing scheme with a scalar parameter $p$, our target is to compute
$\frac{Z(p_{0})}{Z(p)}=\bigg{\langle}\frac{W(p_{0})}{W(p)}\bigg{\rangle}_{p}$ (5)
where $p_{0}$ is the reference point in the parameter space. For the models we consider, $p_{0}=0$. Practically, we take $p_{0}$ to be a sufficiently small number (e.g. $10^{-8}$ in our paper) in the computer. The adequacy can be checked by comparing the results with that of a smaller $p_{0}$ so as to see whether the results converge.
We first take $p=\beta$ for illustration and the case that $p=s$ is similar. The difference will be clarified at the end of this section. Without loss of generality, we assume $\beta_{0}<\beta$. In case that $\beta_{0}$ and $\beta$ are far away in the parameter space, to ensure importance sampling, we divide $[\beta_{0},\beta]$ into $m$ intervals $[\beta_{0},\beta_{1}]$, $[\beta_{1},\beta_{2}]$, $\cdots$, $[\beta_{m-1},\beta_{m}]$, where $\beta_{m}\equiv\beta$, i.e.
Each interval $[\beta_{k-1},\beta_{k}]$ ought to be sufficiently narrow so that the distribution at $\beta_{k}$ can well samples the configurations at $\beta_{k-1}$. Otherwise, the sampling process would be inefficient. Then the natural questions following are how to partition $[\beta_{0},\beta]$ and the how many intervals we need for Eq. (6)
It can be expected that if $\beta_{k}$ and $\beta_{k-1}$ are separated on the two sides of the critical point, we may need a bulk of samples to estimate $Z(\beta_{k-1})/Z(\beta_{k})$. However, we point out that this actually affects little to our algorithm since numerical simulations only deal with finite system, where the criticality does not result in significant inefficiency. We have also verified this point in our simulations. For example, in the 2D spin-$1/2$ XXZ model, we implemented an equal division strategy with step size $\Delta T=0.01$. With $20000$ Monte Carlo samples for each bin, we were able to achieve a sufficiently smooth curve for further numerical differentiation. However, the strategy that we divide $[\beta_{0},\beta]$ into equal intervals or a geometric progression is not economic. We can expect that when $\beta\to 0$ (approaching infinite high temperature), the division can be more sparse compared to that at low temperatures. Therefore, an economic strategy to divide the interval is needed.
In addition to the division problem, $Z(\beta_{0})/Z(\beta)$ could be an exponentially small number exceeding the data range of computer (e.g. $1.7\times 10^{-308}\sim 1.7\times 10^{308}$ for double type data), and we should store $\ln[Z(\beta_{0})/Z(\beta)]$ in the computer. However, simply applying the logarithm to Eq. (6) and converting the multiplication to addition may lead to precision loss in computer calculations if $\ln[Z(\beta_{k-1})/Z(\beta_{k})]$ and $\ln[Z(\beta_{k})/Z(\beta_{k+1})]s$ differ by several orders of magnitude.
To address the two problems above, we present a dynamical approach here. Define $\alpha_{k}\equiv\beta_{k-1}/\beta_{k}<1$, and within the framework of SSE, we have
$\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\langle\alpha^{n_{k}}_{k}\rangle_{\beta_{k}}$ (7)
where $n_{k}$ is the number of all non-identity operators in the SSE simulation. Our dynamical process then starts from the last point with $k=m$, and we require
$\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\epsilon<1,$ (8)
where the choice of $\epsilon$ should match $N_{\text{MC}}$, the number of MC steps we use for samplings, as the error in MC simulations is proportional to $1/\sqrt{N_{\text{MC}}}$.
Next, we set $n_{k}\approx\tilde{n}_{k}=\Lambda_{k}\beta_{k}L$ ($\Lambda_{k}\beta_{k}L^{2}$) for 1D (2D) system, and $\Lambda_{k}$ can be ascertained through the examination of result convergence. This comes from the fact that in SSE, the energy satisfies $E_{k}=-\langle n_{k}\rangle/\beta_{k}$. Practically, we set $\Lambda_{k}$ to be the same for all $k$ for convenience. Then $Z(\beta_{k-1})/Z(\beta_{k})$ can be obtained from the simulation from
$\begin{split}\langle\alpha_{k}^{n_{k}}\rangle=\langle(\epsilon^{1/\tilde{n}_{k}})^{n_{k}}\rangle_{\beta_{k}}\approx\epsilon,\quad\beta_{k-1}=\alpha_{k}\beta_{k}\end{split}$ (9)
Eq. (9) not only requires $Z(\beta_{k-1})/Z(\beta_{k})\approx\epsilon$ for all $k$, but determines the next step in the annealing path. Therefore, with this dynamical annealing strategy, the division will be more sparse when $\beta$ gets smaller.
This strategy actually requires the number divisions of scale proportional to the system size for large systems. For the 1D system as an example,
or
$\beta_{k}-\beta_{k-1}=-\frac{\ln\epsilon}{\Lambda L}$ (11)
then the number of divisions is around $(\beta-\beta_{0})\Lambda L/|\ln\epsilon|$.
One more thing we have to pay attention is that when $\beta\to 0$, we have $\alpha_{k}\to 0$, which will also exceed the range of a number saved in the computer. Therefore we manually set a lower bound for $\alpha_{k}$ (e.g. $\min{\alpha_{k}}=\epsilon$ in our simulation). The existence of lower bound makes the division in the region that $\beta\to 0$) denser than that Eq. (9) requires. Since most $Z(\beta_{k-1})/Z(\beta_{k})$ still at the same order of magnitude, e.g. $\epsilon=10^{-2}$ used for the first two example in our simulation, and the more divisions when $\beta\to 0$ only constitutes a very small proportion, this affects little to the precision. In Fig. 1, we plot the number of divisions in the T-Re-An method of first example in our paper for reference, where we take $\Lambda=L$.
Figure 1: The number of divisions in the T-Re-An method of first example in our paper, where we take $\beta=30$ and $\Lambda=L$.
For the quantum reweight-annealing scheme, we adopt a similar procedure as introduced above. For $Z(s_{k-1})/Z(s_{k})=\langle(s_{k-1}/s_{k})^{n_{k,\text{odd}}}\rangle_{s_{k}}$, we can expect that when $s\to 0$, the number of the operators of the bonds with odd indices goes to zero in Eq. (3). Therefore we can take $n_{k,\text{odd}}\approx\tilde{n}_{k,\text{odd}}=\Gamma_{k}\beta s_{k}L$ ($\Gamma_{k}\beta s_{k}L^{2}$) for 1D (2D) system, and we find this hypothetical linear relation on $s$ is convenient and has sufficiently good performance. The number of divisions in this case similarly scales with $\mathcal{O}(\Gamma\beta L)$ for some uniform $\Gamma_{k}=\Gamma$ if in 1D system.
A.3 (III) Computing entanglement entropy

### v5

@@ III Examples
For the pseudo-annealing scheme, we adopt a similar procedure as in Sec. II.4. For $Z(s_{k-1})/Z(s_{k})=\langle(s_{k-1}/s_{k})^{n_{k,\text{odd}}}\rangle_{s_{k}}$, we can expect that when $s\to 0$, the number of the operators of the bonds with odd indices goes to zero. Therefore we can take $n_{k,\text{odd}}\approx\tilde{n}_{k,\text{odd}}=\Gamma_{k}\beta s_{k}L$ ($\Gamma_{k}\beta s_{k}L^{2}$) for 1D (2D) system, and we find this hypothetical linear relation on $s$ is convenient and has sufficiently good performance just as the first example shows. The number of divisions in this case similarly scales with $\mathcal{O}(\Gamma\beta L)$ for some uniform $\Gamma_{k}=\Gamma$ if in 1D systems.
As it is shown in Fig. 3, both the results of T-Re-An and Q-Re-An method can match that of the exact diagonalization (ED) method exceptionally well for small lengths, which verify the validity of our methods. The results are also consistent with the expectation that $\ln Z\propto L+o(L)$, where $o(L)$ means higher order terms which is negligible compared with $L$. One may feel this example too trivial and easy to be calculated. Therefore we introduce the second example which needs a high-precision to extract the correction of the sub-leading term in entropy.
@@
III.2 2D LBW-HM
To further exploit the advantage of our method, we calculate the thermodynamic entropy of a 2D lattice Bisognano-Wichmann Heisenberg model (LBW-HM) [72, 73], the thermodynamic entropy $S$ of the which at $\beta=1$ provides a good approximation for the Von Neumann entanglement entropy of a 2D PBC square lattice spin-1/2 AFM Heisenberg model under smooth bipartition [74, 75]. A brief review of this theory can be found in Appx. B.
According to the CFT prediction and previous numerical results [76, 77, 57, 78], the coefficient of sub-leading term of the entropy reflects the number of Goldstone modes. Here, the thermal entropy of the LBW-HM at $\beta=1$ should obey $S(L)=aL+b\ln L+c$, and $b=n_{G}/2$, where $n_{G}=2$ is the number of Goldstone modes of the 2D spin-1/2 AFM Heisenberg model [79, 77]. In this context, coarse result of entropy is disfavored because the the coefficient $b$ in front of the sub-leading term is sensitive to errors. We have compared our method with the specific heat integral ($C$-int) and Wang-Landau (WL) algorithm in this example.
@@
For the WL algorithm, the number of density of states to be flattened in each histogram should be greater than $\beta E(\beta)$, where $E(\beta)$ is the energy at $\beta$, which is also a (roughly) polynomial complexity with the system size [15]. However, the main problem of this algorithm is the so-called error saturation, which is intrinsic, preventing us from obtaining results of an arbitrary precision generally and making this algorithm basically biased [80, 81, 18, 19]. In Fig. 4, we show great error of the results obtained from the vanilla WL samplings [15], and to achieve an arbitrary precision within the Wang-Landau framework is not generally easy. Another disadvantage of the WL sampling compared with our method is that it cannot be parallelized: the two subroutines for flattening the corresponding histograms before and after an adjustment of the refinement parameter, must be executed sequentially. In contrast, our unbiased method is able to achieve an arbitrary precision, easy to parallelize, and much more accessible as we only need to measure the observable of $Z(\vec{p}_{k-1})/Z(\vec{p}_{k})$ in many conventional and mature MC algorithms.
III.3 2D XXZ model
As the last example, we consider a much more complex model, a 2D spin-1/2 XXZ model with next nearest neighbor interaction,
@@
Here we use equal division of $T=1/\beta=0.01$, which is convenient for further numerical differentiation. We found that with step size $\Delta T=0.01$ and $20000$ MC samples for each bin, a sufficiently smooth curve can be obtained. Because a constant factor of the PF does not affect the (higher order) derivative of $F$, we choose $Z(T^{*}=1.95)$ as the reference point rather than some known point. The PF we calculated is thus $Z(T)/Z(T^{*})$ where $Z(T^{*})$ is an unknown constant. Then we indeed observe a divergent peak of the second order derivative of the free energy as shown in Fig. 5(a), which accurately probes the phase transition point, matching with the intersection point of Binder ratios $R_{2}=\langle M^{4}\rangle/\langle M^{2}\rangle^{2}$ of different sizes shown in Fig. 5(b), where $M$ denotes the magnetization.
@@ IV Summary and discussions
We propose an unbiased algorithm within the Monte Carlo framework, which has low-technical barrier and no systemic error. The method can be easily implemented for parallel computation and can obtain the value of partition function, free energy and thermal entropy with high precision. The method has two keynotes: (1) Finding a parameter path in whole space (either temperature or external parameter) connecting a known point and the to-be-solved one. If we only want to calculate the (high order) derivative of the PF, even fixing an unknown point as the reference is enough; (2) Annealing along the path to calculate the final result through a propagation formula.
@@
We show the advantage of our method via calculating the thermal entropy of 2D LBW-HM and successfully extract precise coefficient of the sub-leading term ($\ln L$), which is consistent with the CFT prediction and catch the correct number of Goldstone modes. In comparison, the $C$-int and the WL sampling both have worse performance. In a 2D XXZ model, the numerical second derivative of the Re-An data is still smooth and precise, which accurately probes the phase transition point. Our method makes the efficient and high-precision calculation of entropy and other related physical quantities accessible, which can reveal some universal information, such as CFT, criticality and symmetry breaking.
The final thing worth of emphasizing is that the reweight-annealing method discussed here is a general paradigm, not limited in computing the standard partition function, but can be extended to any generalized partition function as long as the two keynotes can be satisfied. One representative example has been shown in [82], where the authors developed an algorithm to compute the Rényi entropy using this reweight-annealing framework.
Acknowledgement
@@
We acknowledge the start-up funding of Westlake University. The authors also acknowledge Beijing PARATERA Tech Co.,Ltd.(https://www.paratera.com/) for providing HPC resources that have contributed to the research results reported within this paper. CC and ZY thank the support of the National Natural Science Foundation of China (grant no. 12247101). NM thanks the support of the National Natural Science Foundation of China (grant no. 12004020). GP thanks the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter - ct.qmat (EXC 2147, Project No. 390858490).
Note: After we finished this work, we found a related work posted on arXiv two weeks ago [83].
@@
Appendix A Estimators in thermal and quantum reweight-annealing under the SSE framework
We first review the reweighting trick in a general MC method with the inverse temperature $\beta$ for illustrations (the thermal case), then we will return to the SSE method, which is just one way of transforming the quantum degrees of freedom to classical ones in the partition function.
Consider a partition function $Z(\beta)=\sum_{C}W(\beta;C)$, where the weights $\{W(\beta;C)\}$ are non-negative real numbers (such that we can perform MC simulations) and $C$ denotes each configuration. For the expectation value of some diagonal operator $O$ at the inverse temperature $\beta^{\prime\prime}$, This can be achieved as follows:
$\begin{split}\langle O\rangle_{\beta^{\prime\prime}}=&\frac{\sum_{C}W(\beta^{\prime\prime};C)O(C)}{Z(\beta^{\prime\prime})}\\ \end{split}$ (11)
where $O(C)$ is the value of the operator $O$ under the configuration $C$. In our paper, a specific form of the operator $O=[W(\beta^{\prime})/W(\beta^{\prime\prime})]$ is discussed, then we have
@@ Appendix A Estimators in thermal and quantum reweight-annealing under the SSE framework
$\begin{split}\bigg{\langle}\frac{W(\beta^{\prime})}{W(\beta^{\prime\prime})}\bigg{\rangle}_{\beta^{\prime\prime}}=&\frac{1}{Z(\beta^{\prime\prime})}\sum_{C}W(\beta^{\prime\prime};C)\frac{W(\beta^{\prime};C)}{W(\beta^{\prime\prime};C)}\\ =&\frac{\sum_{C}W(\beta^{\prime};C)}{Z(\beta^{\prime\prime})}\\ =&\frac{Z(\beta^{\prime})}{Z(\beta^{\prime\prime})}\end{split}$ (12)
Eq. (12) is the key relation in this paper, in which we sample the configurations with probabilities $p(\beta^{\prime\prime};C)=W(\beta^{\prime\prime};C)/Z(\beta^{\prime\prime})$ at inverse temperature $\beta^{\prime\prime}$, but recalculating the weights of the sampled configuration $C$ at another inverse temperature $\beta^{\prime}$, i.e., the $W(\beta^{\prime};C)$. What Eq. (12) indicates is that through averaging the weight ratio of each configuration, the ratio of partition functions at different inverse temperature $Z(\beta^{\prime})/Z(\beta^{\prime\prime})$ can be obtained.
Now we turn to SSE, the demonstrative QMC method used in this paper [1, 2]. Apply the Taylor expansion, we have
$Z(\beta)=\text{tr}\bigg{\{}\sum_{n}\frac{\beta^{n}}{n!}(-H)^{n}\bigg{\}}=\sum_{n}\frac{\beta^{n}}{n!}\text{tr}\{(-H)^{n}\}$ (13)
Then we write $H=-\sum_{a,b}H_{a,b}$, where $H_{a,b}$ are some non-identity local operator with $a$ denoting the type of it (diagonal or off-diagonal under the simulating basis) and $b$ denotes the index of the space degree (bond or site) that the local operator is acting on. Hence
$(-H)^{n}=\sum_{S_{n}}\prod_{p=1}^{n}H_{a_{p},b_{p}}$ (14)
where $S_{n}$ denotes that we sum over all possible operator strings, and each of them is a product of $n$ local operators. Then we have
$Z(\beta)=\sum_{n}\sum_{S_{n}}\frac{\beta^{n}}{n!}\text{tr}\bigg{\{}\prod_{p=1}^{n}H_{a_{p},b_{p}}\bigg{\}}$ (15)
The important thing here is that because the product of $n$ local operators are interpreted as an operator string, therefore each string $S_{n}$ has exactly $n$ non-identity operators. In other words, each $S_{n}$ has the length $n$.
Replacing the trace operation with a set of complete basis $\{\alpha_{0}\}$, we have
$Z(\beta)=\sum_{n}\sum_{S_{n}}\sum_{\alpha_{0}}\frac{\beta^{n}}{n!}\bigg{\langle}\alpha_{0}\bigg{|}\prod_{p=1}^{n}H_{a_{p},b_{p}}\bigg{|}\alpha_{0}\bigg{\rangle}$ (16)
Since practically we cannot simulate $n$ to infinity, we introduce a large cut-off $M$ to truncate the series. For those $n<M$, after inserting some identity operators (also some special $H_{a,b}$ by using $a$ to denote identity, diagonal and off-diagonal types), we can make all operator strings have the same length $M$. Besides, we introduce $(M-1)$ extra sets of complete basis, and finally we have
$Z(\beta)=\sum_{S_{M}}\sum_{\alpha}\frac{\beta^{n}(M-n)!}{M!}\prod_{p=1}^{M}\langle\alpha_{p-1}|H_{a_{p},b_{p}}|\alpha_{p}\rangle$ (17)
where we have identify $\alpha_{M}\equiv\alpha_{0}$ and $\alpha\equiv\alpha_{0},\cdots,\alpha_{M-1}$.
The weight in Eq.(17) is
$W(\beta;S_{M},\alpha)=\frac{\beta^{n}(M-n)!}{M!}\prod_{p=1}^{M}\langle\alpha_{p-1}|H_{a_{p},b_{p}}|\alpha_{p}\rangle$ (18)
and the classical configurations are given by both the states $\alpha$ and the operator string $S_{M}$. By replacing the weights $W(\beta^{\prime};C)$ and $W(\beta^{\prime\prime};C)$ in Eq. (12) with $W(\beta^{\prime};S_{M},\alpha)$ and $W(\beta^{\prime\prime};S_{M},\alpha)$, we have
$\frac{Z(\beta^{\prime})}{Z(\beta^{\prime\prime})}=\bigg{\langle}\bigg{(}\frac{\beta^{\prime}}{\beta^{\prime\prime}}\bigg{)}^{n}\bigg{\rangle}_{\beta^{\prime\prime}}$ (19)
for $n$ to denote the number of non-identity operators.
@@
Similar discussions can be applied to the quantum case. For $H(s)=sH_{0}+(1-s)H_{1}=H_{1}+s(H_{0}-H_{1})$, we usually have $Z(s^{\prime})/Z(s^{\prime\prime})=\langle(s^{\prime}/s^{\prime\prime})^{n_{s}}\rangle_{s^{\prime\prime}}$ within the framework of SSE, where $n_{s}$ is the number of non-identity operators related to $(H_{0}-H_{1})$. For instance, if
@@
where $H_{0}=\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$ and $H_{1}=\sum_{i\in\text{even}}\vec{S}_{i}\cdot\vec{S}_{i+1}$, then we can obtain
@@
where $n_{\text{odd}}$ ($n_{\text{even}}$) denotes the amounts of non-identity operators acting on the bonds with odd (even) indices in SSE simulations, and
$\tilde{S}_{\alpha,M}=\frac{\beta^{n}(M-n)!}{M!}\prod_{p=1}^{M}\langle\alpha_{p-1}|H_{a_{p},b_{p}}|\alpha_{p}\rangle.$ (22)
Appendix B Computing entanglement entropy